letae R and Ma = {x,y) ER?: y2 = x(x - 1)(x - a)} for which values of a is M, an embedded sub manifold of R2?
Answer #1

[Son] so, let qel and Ma= {ley ER² y = 11-1) (2-a) Falow y) = y24127960-a) Ma = if (0) Jelisey = {l-34291 +232 – 9,247 (x, y)